OpenStudy (anonymous):

In a group of 20 singers and 40 dancers, 20% of the singers are under 25 years old and 40% of the entire group are under 25 years old. What percent of he dancers are under 25 years old? 1. 20% 2. 40% 3. 50% 4. 60% 5. 80%

OpenStudy (anonymous):

20% of 20 singers r under 25! dat makes it 20/100x20=4 singers.40%of 60 ppl r under 25,d no. tirns out 2 b 40/100 x 60=2/5 x 60=24 .dis implies dat there r 20 dancers under d age ov 25!20/40 x 100=50%

OpenStudy (anonymous):

thankie :)

OpenStudy (anonymous):

yw!!